### v1

@@ 5.1 Experimental Setting
 Sphere  10  1.00  1787  1.00 1146  1.00  1443  1.00  1633
 20  1.00  3329  1.00 2390  1.00  2985  1.00  3249
 40  1.00  6276  1.00 5379  1.00  6020  1.00  6234
 80  1.00  11568  1.00  11995  1.00 11535  1.00  11534
 Ellipsoid  10  1.00  6078  0.11  503818  1.00  6036  1.00 5967
 20  1.00  18906  1.00  26085  1.00  18790  1.00 18773
 40  1.00  68401  1.00  74922  1.00 67119  1.00  68031
 80  1.00  265855  1.00  275238  1.00 260095  1.00  265113
 Cigar  10  1.00  4423  0.58  42639  1.00 3896  1.00  4181
 20  1.00  8693  1.00  8125  1.00 8022  1.00  8514
 40  1.00  16976  1.00 15328  1.00  16418  1.00  16901
 80  1.00  31951  1.00  32579  1.00 31738  1.00  31870
 Rosenbrock  10  0.93  6972  0.37  115873  0.91 6725  0.89  6789
 20  0.90  24015  0.78  37830  0.83 22415  0.90  22870
 40  0.93  87489  0.81  86794  0.83 83574  0.86  91485
 80  0.91  354686  0.79 299309  0.89  306423  0.85  374382
 Ackley  10  0.97  3719  0.71  3323  0.92 3190  0.96  3508
 20  0.95  6946  0.49  10209  0.94 6366  0.99  6512
 40  1.00  12143  0.39  27548  0.98 12038  1.00  12101
 80  1.00  21974  0.00  -  0.97  22802  1.00 21915
 Rastrigin  10  0.99  50781  0.99  50445  0.98  50241  1.00 49931
 20  1.00  167412  0.99  168683  1.00  167174  1.00 166502
 40  1.00 606858  0.98  621647  1.00  608979  0.99  614402
 80  1.00  2110892  0.99  2125240  1.00 2091376  1.00  2091908
@@ 5.2 Results and Discussion
Table 2 shows the success rate and SP1 [5] over 100 independent trials. We note that the SP1 index is defined as the average evaluation counts in successful trials divided by the success rate. The MAP-CMA with $r=1$ suffers from instability except for Sphere. This is due to that the effect of the momentum update, i.e., the old samples accumulated in the evolution path, is too large. In most cases, the MAP-CMA with $r=\sqrt{N}$ is marginally better than the CMA-ES on the SP1 index. Figure 2 shows the transitions of the best evaluation value on the 100 independent trials. For Sphere, Cigar, Rosenbrock, and Ackley, the MAP-CMA improves the best evaluation values more efficiently than the CMA-ES. This is probably because that the momentum update, which contains the evolution path in the mean vector update, reduces the variance of the update direction.
Figure 3 shows the transition of the mean vector when optimizing Rosenbrock with $N=20$. The initial mean vector and step-size were given as $\boldsymbol{m}^{(0)}=(1,\ldots,1)$ and $\sigma^{(0)}=1$, respectively. Once the mean vector reaches near the origin, it moves to $(1,\ldots,1)$. At this stage, the momentum update in the mean vector update allows the MAP-CMA to move the mean vector faster than the CMA-ES.
For Ellipsoid optimization, the performance difference between the CMA-ES and MAP-CMA is small because covariance matrix adaptation is the bottleneck of the optimization. When optimizing Ackley with the MAP-CMA, the optimization stalls in some trials. This stagnation may be due to the effect of past samples [22], which is an issue to be investigated in the future. While the MAP-CMA is mostly successful in Rastrigin, its advantage over the CMA-ES diminishes due to the effect of past samples, similar to the situation observed in Ackley.
@@ 6 Conclusion
We provided a new interpretation of the rank-one update in the CMA-ES from the perspective of the natural gradient with prior distribution. We firstly proposed maximum a posteriori IGO (MAP-IGO), which is the IGO framework extended to incorporate a prior distribution, and derived the update rules of the MAP-IGO in which the multivariate normal distribution and the normal-inverse-Wishart distribution are applied. Then, we derived the rank-one update from the MAP-IGO by setting the prior distribution based on the idea that the promising mean vector should exist in the direction of the evolution path. Furthermore, the newly derived rank-one update is extensible, and we proposed maximum a posterior CMA-ES (MAP-CMA), a CMA-ES equipped with it. The mean update rule in the MAP-CMA has an additional term containing the evolution path, which can be interpreted as a momentum term as discussed in [18] and we call this additional term momentum update in this study. Experimental results showed that the MAP-CMA is marginally better than the CMA-ES in optimizing the various benchmark functions.

### v2

@@ 5.1 Experimental Setting
 Sphere  10  1.00 1787  1.00  2399  1.00  1961  1.00  1824
 20  1.00 3329  1.00  5034  1.00  3617  1.00  3366
 40  1.00 6276  1.00  10719  1.00  6705  1.00  6322
 80  1.00 11568  1.00  24518  1.00  12246  1.00  11658
 Ellipsoid  10  1.00  6078  1.00  6888  1.00  6215  1.00 6050
 20  1.00  18906  1.00  21486  1.00  18984  1.00 18853
 40  1.00  68401  1.00  74256  1.00 67205  1.00  68056
 80  1.00  265855  1.00  284614  1.00 259434  1.00  264651
 Cigar  10  1.00 4423  1.00  5623  1.00  4768  1.00  4506
 20  1.00 8693  1.00  12314  1.00  9256  1.00  8787
 40  1.00 16976  1.00  26003  1.00  17738  1.00  17068
 80  1.00 31951  1.00  58000  1.00  33227  1.00  32098
 Rosenbrock  10  0.93  6972  0.91  7039  0.91  6825  0.94 6802
 20  0.90  24015  0.80  24152  0.92 20599  0.91  22824
 40  0.93  87489  0.76  80442  0.89 77235  0.94  82818
 80  0.91  354686  0.80 258692  0.92   287489  0.91  344089
 Ackley  10  0.97 3719  0.73  6734  0.94  4248  0.96  3964
 20  0.95 6946  0.60  17122  0.96  7555  0.96  6991
 40  1.00 12143  0.31  69112  0.98  13318  1.00  12316
 80  1.00 21974  0.00  -  0.91  25775  0.99  22314
 Rastrigin  10  0.99  50781  1.00 50113  0.98  50722  0.97  51602
 20  1.00 167412  1.00  168168  1.00  167748  1.00  168196
 40  1.00 606858  0.99  620722  1.00  610302  1.00  608874
 80  1.00  2110892  1.00  2127692  1.00 2096612  1.00  2106244
@@ 5.2 Results and Discussion
Table 2 shows the success rate and SP1 [5] over 100 independent trials. We note that the SP1 index is defined as the average evaluation counts in successful trials divided by the success rate. The behavior of the MAP-CMA approaches that of the CMA-ES as $r$ increases. When $r$ is too small and the effect of the old samples accumulated in the evolution path is too large, the momentum update in the MAP-CMA tends to have a negative effect on optimization. When $r=N$, the MAP-CMA is superior to the CMA-ES for Rosenbrock and competitive for the other functions. Figure 2 shows the transitions of the best evaluation value on the 100 independent trials. While the difference in performance between the CMA-ES and the MAP-CMA is slight for Sphere, Ellipsoid, Cigar, and Rastrigin, the MAP-CMA improves the best evaluation values more efficiently than the CMA-ES for Rosenbrock.
Figure 3 shows the transition of the mean vector when optimizing Rosenbrock with $N=20$. The initial mean vector and step-size were given as $\boldsymbol{m}^{(0)}=(1,\ldots,1)$ and $\sigma^{(0)}=1$, respectively. Once the mean vector reaches near the origin, it moves to $(1,\ldots,1)$. At this stage, the momentum update in the mean vector update allows the MAP-CMA to move the mean vector faster than the CMA-ES. This result suggests that the MAP-CMA can be more efficient than the CMA-ES when the required mean vector moves are large.
When optimizing Ackley with the MAP-CMA, the optimization stalls in some trials. This stagnation may be due to the effect of past samples [22], which is an issue to be investigated in the future.
@@ 6 Conclusion
We provided a new interpretation of the rank-one update in the CMA-ES from the perspective of the natural gradient with prior distribution. We firstly proposed maximum a posteriori IGO (MAP-IGO), which is the IGO framework extended to incorporate a prior distribution, and derived the update rules of the MAP-IGO in which the multivariate normal distribution and the normal-inverse-Wishart distribution are applied. Then, we derived the rank-one update from the MAP-IGO by setting the prior distribution based on the idea that the promising mean vector should exist in the direction of the evolution path. Furthermore, the newly derived rank-one update is extensible, and we proposed maximum a posterior CMA-ES (MAP-CMA), a CMA-ES equipped with it. The mean update rule in the MAP-CMA has an additional term containing the evolution path, which can be interpreted as a momentum term as discussed in [18] and we call this additional term momentum update in this study. Experimental results showed that the MAP-CMA is better than the CMA-ES in optimizing functions that require a lot of mean vector moves, such as Rosenbrock.